# Non-Holomorphic $A_{4}$ Modular Symmetry in Type-I Seesaw: Implications for Neutrino Masses and Leptogenesis

### v1

@@
We propose a minimal extension of the Standard Model with right-handed neutrinos, governed by a non-holomorphic $A_{4}$ modular flavor symmetry of level $N=3$. Within this model framework, the light neutrino masses are generated via the popular type-I seesaw mechanism in which the structure of the Dirac neutrino Yukawa couplings is decided by nonholomorphic modular forms. Unlike conventional flavor models with ad hoc flavon fields, the structure of Dirac and Majorana mass matrices is entirely determined by a modulus parameter $\tau$. We construct the predictive mass matrices for charged leptons, Dirac neutrinos, and right-handed Majorana neutrinos and show the compatibility with neutrino oscillation data by an appropriate choice of input model parameters. We find that our $\chi^{2}$ analysis of neutrino masses and mixing gives excellent agreement with current neutrino oscillation observables by taking normal hierarchical pattern of light neutrinos. We present numerical analysis of two sets of benchmark points explaining neutrino masses while generating the correct amount of baryon asymmetry via thermal leptogenesis. We estimate numerically the values of CP-asymmetry and examine the evolution of the lepton asymmetry by studying Boltzman equations by considering both strong and washout regimes with CP-asymmetry parameter in the range $|\varepsilon_{1}|\sim 10^{-4}$$10^{-8}$. The model predicts an effective Majorana mass in the few meV range, below current experimental bounds but within reach of next-generation $0\nu\beta\beta$ searches. The key feature of non-holomorphic $A_{4}$ modular symmetry naturally accommodates non-zero neutrino masses and mixings, minimizes the Yukawa arbitrariness, and establishes a direct connection between high-scale leptogenesis with low-energy neutrino observable parameters, thereby the model provides a testable link between neutrino flavor physics and cosmology.
@@ 1 Introduction
Even after these spectacular findings, there are some fundamental questions at the the theoretical and experimental level that remained to be answered in order to understand neutrino physics. The well-determined values of neutrino mass-square differences and mixing angles motivated to propose a simpler extension of the SM to understand the origin and structure of neutrino masses and mixing. One of the elegant idea to explain non-zero neutrino masses and mixing is popularly known as seesaw mechanism, where heavy messenger degree of freedom couple to SM neutrino, can generate small masses for light active neutrinos naturally through a suppression factor with a large messenger mass scale. The most minimal and well-studied realization is the type-I seesaw mechanism [10, 11, 12, 13], which introduces three SM-singlet right-handed neutrinos, allowing the light neutrino masses to emerge as the low-energy remnants of a high-scale Majorana mass matrix. Alternative seesaw mechanisms, including type-II [14, 15, 16, 17, 18] and type-III [19, 20], introduce scalar triplets and fermion triplets respectively, offering different model-building avenues.
@@
The structure of the paper is as follows. We present the adequate theoretical formulation of modular $A_{4}$ symmetry including transformation properties of particles and Yukawa couplings and assignment of proper modular weight in Sec-2, In Sec-3, we discuss the methodology of detailed numerical analysis and parameter scanning. We perform the $\chi^{2}$ analysis and show the phenomenological results for neutrino oscillation observables in Sect.4, we discuss the role of modular $A_{4}$ symmetry in estimating CP-asymmetry and study the evolution of lepton asymmetry via the Boltzmann equations to explain correct baryon asymmetry of the universe via themral leptogensis. Finally, we conclude the summary of our findings in Sect.5.
@@ 3.1 Structure of the charged lepton mass matrix
where $Y^{(0)}_{3}$ denotes the $A_{4}$ modular triplet constructed from polyharmonic Maass forms (PHMFs) of weight zero. The need for flavon fields are replaced by these modular forms and contribute to the role of effective Yukawa couplings. Their explicit Fourier expansions are provided in Eqns (B), (B).
@@ 4 Numerical Analysis of Neutrino Masses and Mixings
To begin the numerical analysis, we chose the values of ${\rm Re}[\tau]$ and ${\rm Im}[\tau]$ over a range of $|{\rm Re}[\tau]|\in$[0, 0.5] and ${\rm Im}[\tau]\in$ [0.9, 1.2] such that it satisfies the constraints of the most generic fundamental domain i.e, $|{\rm Re}[\tau]|\leq 1/2$, ${\rm Im}[\tau]>0$ and $|\tau|\geq 1$. The Yukawa couplings used in our model are evaluated using the values of ${\rm Re}[\tau]$ and Im($\tau$) from their q-expansions as given in Eqns (B), (B). We also fix phenomenologically motivated other model parameters like $|\alpha_{D},|\;|\beta_{D}|,\;|\gamma_{D}|$ over a range of [0.1, 0.99] and $\Lambda_{1},\;\Lambda_{2}$ for a range of $\Lambda_{1}\in[1.0\times 10^{13},9\times 10^{13}]$ GeV and $\Lambda_{2}\;\in\;[1.0\times 10^{14},9\times 10^{14}]$ GeV, respectively. These input values are then used to find the numerical values of $M_{D}$, $M_{R}$ and eigenvalues of the neutrino mass matrix (Eqn. 3.4).
@@
$\theta_{12}(^{\circ})$ $\theta_{23}(^{\circ})$ $\theta_{13}(^{\circ})$ $\Delta m^{2}_{21}/10^{-5}(\text{eV}^{2})$ $\Delta m^{2}_{31}/10^{-3}(\text{eV}^{2})$
33.18 48.66 8.31 7.79 2.48
$\delta_{CP}(^{\circ})$ $m_{e}$ (MeV) $m_{\mu}$ (MeV) $m_{\tau}$ (MeV) $\tau$
176.58 0.51 105.65 1776.96 0.216 + 0.996 i
$m_{1}$ (eV) $m_{2}$ (eV) $m_{3}$ (eV)
0.0016 0.0089 0.0498
$\alpha(^{\circ})$ $\beta(^{\circ})$ $\Lambda_{1}$ (GeV) $\Lambda_{2}$ (GeV) $Y^{(0)}_{3,1}$
88.94 101.73 $3.70\times 10^{12}$ $2.35\times 10^{13}$ $0.209+4.33\times 10^{-19}i$
$Y^{(0)}_{3,2}$ $Y^{(0)}_{3,3}$ $Y^{(-2)}_{3,1}$ $Y^{(-2)}_{3,2}$ $Y^{(-2)}_{3,3}$
$0.179+0.052i$ 0.1790.052i0.179-0.052i 0.420.0145i0.42-0.0145i 0.0946+0.090i-0.0946+0.090i 0.253+0.128i-0.253+0.128i
$|\alpha_{l}/\beta_{l}|$ $|\beta_{l}/\gamma_{l}|$ $|\alpha_{D}|$ $|\beta_{D}|$ $|\gamma_{D}|$
0.6365 1.198 0.1074 0.1229 0.073
Table 4: The above table presents the best-fit values for BP1 with Low $M_{D}$. The $\chi^{2}_{min}\approx 0.000267$ for this set of measurements. The benchmark point for this is already presented in Table [5]. The best-fit values of the observables and the fitted parameters shown above corresponds to the minimum $\chi^{2}$ value that we have obtained for Normal Hierarchy.
@@
Table 5: Representative benchmark points obtained from $\chi^{2}$ fit of the model parameters with different choice of derived $M_{D}$ and fixed value of $M_{R}$. The parameters used for numerical analysis includes BP1 (2nd column) with Low $M_{D}$, $\Lambda_{1}\approx\mathcal{O}(10^{12}\;\mbox{GeV})\;\text{and}\;\;\Lambda_{2}\approx\mathcal{O}(10^{13}\;\mbox{GeV})$ and BP2 (3rd column) with High $M_{D}$, $\Lambda_{1}\approx\mathcal{O}(10^{13}\;\mbox{GeV})\;\text{and}\;\;\Lambda_{2}\approx\mathcal{O}(10^{14}\;\mbox{GeV})$.
4.1 $\chi^{2}$ analysis and benchmark point
@@
$\chi^{2}=\sum_{n}\left(\dfrac{\mbox{P}^{\rm th}_{\rm n}-\mbox{P}^{\rm exp}_{\rm n}}{\sigma_{n}}\right)^{2}$ (4.5)
The set of observables included in the sum are neutrino parameters $\{\,\theta_{21}$, $\theta_{31}$, $\theta_{32}$, $\Delta m^{2}_{21}$$\Delta m^{2}_{31}$, $\delta_{CP}\,\}$ and mass ratios for charged leptons $m_{e}/m_{\mu}$, $m_{\mu}/m_{\tau}$. In our case, we simply follow the standard $\chi^{2}$ minimization procedures  [^1] One can also perform this similar simulation on the $\chi^{2}$ minimization either by using the MultiNest sampling package [115] or the software package MINUIT [116]. using Eqn (4.5) and present the best fit of the model parameters corresponding to $\chi^{2}_{min}$ in Table [4] (named BP1) and Table [6] (named as BP2). These two representative benchmark points (BPs) are selected based on two different structures of Dirac neutrino mass matrix to show its impact on the high scale phenomenology like thermal leptogenesis.
$\theta_{12}(^{\circ})$ $\theta_{23}(^{\circ})$ $\theta_{13}(^{\circ})$ $\Delta m^{2}_{21}/10^{-5}(\text{eV}^{2})$ $\Delta m^{2}_{31}/10^{-3}(\text{eV}^{2})$
33.46 48.77 8.24 7.98 2.60
$\delta_{CP}(^{\circ})$ $m_{e}$ (MeV) $m_{\mu}$ (MeV) $m_{\tau}$ (MeV) $\tau$
176.86 0.510 105.65 1776.87 0.216 + 0.997 i
$m_{1}$ (eV) $m_{2}$ (eV) $m_{3}$ (eV)
0.0016 0.0090 0.0510
$\alpha(^{\circ})$ $\beta(^{\circ})$ $\Lambda_{1}$ (GeV) $\Lambda_{2}$ (GeV) $Y^{(0)}_{3,1}$
309.45 28.25 $8.69\times 10^{13}$ $6.63\times 10^{14}$ $0.209+4.33\times 10^{-19}i$
$Y^{(0)}_{3,2}$ $Y^{(0)}_{3,3}$ $Y^{(-2)}_{3,1}$ $Y^{(-2)}_{3,2}$ $Y^{(-2)}_{3,3}$
$0.18+0.052i$ 0.180.052i0.18-0.052i 0.4200.015i0.420-0.015i 0.098+0.0947i-0.098+0.0947i 0.257+0.130i-0.257+0.130i
$|\alpha_{l}/\beta_{l}|$ $|\beta_{l}/\gamma_{l}|$ $|\alpha_{D}|$ $|\beta_{D}|$ $|\gamma_{D}|$
1.7155 0.6750 0.5228 0.4336 0.5811
Table 6: The above table presents the best-fit values for BP2 with High $M_{D}$. The $\chi^{2}_{\rm min}\approx 0.000169$ for this set of measurements. The benchmark point for this is already presented in Table 5. The best-fit values of the observables and the fitted parameters shown above corresponds to the minimum $\chi^{2}$ value that we have obtained for Normal Hierarchy.
[…]
(a) Re[$Y^{(-2)}_{3,2}$] vs Re[$\tau$]
(b) Re[$\tau$] vs Im[$\tau$]
(c) $\chi^{2}$ scale
Figure 1: Correlation plots for benchmark point BP1: (a) Re[$Y^{(-2)}_{3,2}$] vs Re[$\tau$], (b) Re[$\tau$] vs Im[$\tau$], and (c) vertical colour bar showing the $\chi^{2}$ values used for color coding of points in (a)–(b). Blue (red) points correspond to lower (higher) $\chi^{2}$.
(a) Re[$Y^{(-2)}_{3,2}$] vs Re[$\tau$]
(b) Re[$\tau$] vs Im[$\tau$]
(c) $\chi^{2}$ scale
Figure 2: Correlation plots for benchmark point BP2: (a) Re[$Y^{(-2)}_{3,2}$] vs Re[$\tau$], (b) Re[$\tau$] vs Im[$\tau$], and (c) vertical colour bar showing the $\chi^{2}$ values used for color coding of points in (a)–(b). Blue (red) points correspond to lower (higher) $\chi^{2}$.
@@
(a) $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{13}$
(b) $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{23}$
(c) $\chi^{2}$ scale
Figure 3: Correlation plots between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ (in Fig.3(a)) and between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{23}$ (in Fig.3(b)) for benchmark point (BP1) presented in Table.4. The vertical colorbar indicates the variation of $\chi^{2}$ values with blue and red colors corresponds to minimum and maximum $\chi^{2}$ values. The data points in both the correlation points presented above are color-coded according to the value of the $\chi^{2}$ for each point.
@@ 4.2 Correlation plots and modular dependence
We have presented the correlation plots between the real part of $Y^{(-2)}_{3,2}$ and the real part of the modulus parameter $\tau$ in Fig.1(a) as well as the correlations among real and imaginary parts of $\tau$ is shown in Fig.1(b), within the parameter space allowed $3\sigma$. These correlations as shown in Fig.1(a) and Fig.1(b) for BP1 and in Fig.2(a) and Fig.2(b) in BP2 illustrate that the modular parameter $\tau$ mainly controls the component of modular Yukawa triplets $Y^{(-2)}_{3,i}(\tau)$. Any small change in ${\rm Re}[\tau]$ or ${\rm Im}[\tau]$ results in smooth changes in triplet components. The allowed model parameters form narrow bands, demonstrating that the allowed value of $\tau$ fixed by the neutrino observables determines the modular Yukawa triplets.
@@
(a) $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{13}$
(b) $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{23}$
(c) $\chi^{2}$ scale
Figure 4: Plot showing correlations between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ (in Fig.4(a)) and between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{23}$ (in Fig.4(b)) for benchmark point (BP2) presented in Table.6. The vertical colorbar indicates the variation of $\chi^{2}$ values as shown in Fig.4(c). The blue and red colors corresponds to minimum and maximum $\chi^{2}$ values.
Furthermore, Fig. 3(a) and Fig. 3(b) in BP1 and Fig. 4(a) and Fig. 4(b) in BP2 show the correlation plots of the mixing angles within the allowed parameter space $3\sigma$. This analysis shows that our model exhibits strong predictability of the neutrino oscillation parameters that agree with the current experimental limits of the mixing angles. These correlations exhibit an elliptical contour rather than a uniform distribution, which is a direct consequence of a reduced number of model parameters because of a modular type-I seesaw framework. Another interesting prediction of these angle-angle correlation plots is the low $\chi^{2}$ regions in both Figs. 3-4 form narrow loci in the angle-angle plane. Thus, the mixing angles computed in BP1 and BP2 agree with the NuFit values and prefer normal ordering. The correlations plot in the planes of light neutrino masses and mass square difference parameters are shown in Fig. 5(a) and Fig. 5(b) for BP1 for $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{13}$ and in Fig. 6(a) and Fig. 6(b) for BP2 for $\sin^{2}\theta_{12}$ vs $\sin^{2}\theta_{23}$. The resulting correlations demonstrate that the light neutrino masses $m_{i}$ for normal ordering are not entirely arbitrary rather strongly constrained. For both the above mentioned benchmark points, the predicted value of the lightest neutrino mass is of the order of $\mathcal{O}(\text{meV})$ and the element of the effective neutrino mass $|m_{ee}|$ is predicted to be small (around a few meV), which is well below the current experimental bounds but potentially testable by next-generation and beyond ton-scale experiments.
For the Figs.3-6, we traced the confidence ellipses that give the appropriate $\chi^{2}$ quantiles for our sampled population in scattered plots at the desired confidence levels. The red, purple and green ellipses represent the upper limit of the $1\sigma,\;2\sigma\;\text{and}\;3\sigma$ confidence intervals, i.e., $68\%$, $95\%$ and $99.7\%$ coverage of the population space, respectively. The color of the data points corresponds to the $\chi^{2}$ value obtained for that particular point. The blue color indicates a lower value of $\chi^{2}$, while points with yellow-orange-red indicate higher values of $\chi^{2}$. A lower value of $\chi^{2}$ indicates that the parameter values predicted by our model have not deviated significantly from the experimental best-fit values at the confidence level $1\sigma$. The best-fit values for different choices of $M_{D}$ are presented in Table 4 and Table 6, respectively. In both cases, our predicted value of $\delta_{CP}$ falls within the range of parameter parameters of $3\sigma$. Also, it should be noted that no data points were generated for the inverted hierarchy, which satisfies the constraints of mass-squared differences.
(a) Correlation: $m_{1}$ (eV) vs $m_{2}$ (eV)
(b) Correlation: $m_{1}$ (eV) vs $m_{3}$ (eV)
(c) Correlation: $\Delta m^{2}_{21}$ (eV2) vs $\Delta m^{2}_{31}$ (eV2)
(d) $\chi^{2}$ scale
Figure 5: Correlation plots in the plane of $m_{1}$ and $m_{2}$ in Fig.5(a), $m_{1}$ and $m_{3}$ plane in Fig.5(b) and between $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$ in Fig.5(c), for BP1. The right-most vertical panel is a rotated color bar indicating the lower $\chi^{2}$ values in blue color and higher $\chi^{2}$ values in red color.
(a) Correlation: $m_{1}$ (eV) vs $m_{2}$ (eV)
(b) Correlation: $m_{1}$ (eV) vs $m_{3}$ (eV)
(c) Correlation: $\Delta m^{2}_{21}$ (eV2) vs $\Delta m^{2}_{31}$ (eV2)
(d) $\chi^{2}$ scale
Figure 6: Correlation plots in the plane of $m_{1}$ and $m_{2}$ in Fig.6(a), $m_{1}$ and $m_{3}$ plane in Fig.6(b) and between $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$ in Fig.6(c), for BP2. The right-most vertical panel is a rotated colour bar indicating the lower $\chi^{2}$ values in blue color and higher $\chi^{2}$ values in red color.
@@
Experiment Isotope $T^{0\nu}_{1/2}\;[10^{25}\;yrs]$ $m_{ee}\;(meV)$ Ref
GERDA+MJD+L200 $^{76}Ge$ 1919 7520075-200 [117]
KamLAND-Zen 800 $^{136}Xe$ $38$ 2812228-122 [118]
CUORE $^{130}Te$ $3.8$ 7024070-240 [119]
@@
$T^{0\nu}_{1/2}$ $m_{ee}\;(meV)$
CUPID $Li_{2}^{100}MoO_{4}\;(crystal)$ $1.1\times 10^{27}\;yrs$ 122012-20 [120]
LEGEND-1000 $^{76}Ge$ $1.3\times 10^{28}\;yrs$ 9219-21 [121]
nEXO $^{136}Xe$ $7.4\times 10^{27}\;yrs$ 6276-27 [122]
Future Ultimate Experiments of $0\nu\beta\beta$ decay
@@
$T^{0\nu}_{1/2}$ $m_{ee}\;(meV)$
LEGEND-6000 $10^{29}\;yrs$ 363-6 [123]
ORIGIN-X $10^{30}\;yrs$ <2<2 [124]
THEIA $1.1(2.0)\times 10^{28}\;yrs$ <6.3(5.6)<6.3(5.6) [125]
$90\%\;C.L.\;for\;Te(Xe)$

### v2

@@
We propose a minimal extension of the Standard Model with right-handed neutrinos, governed by a non-holomorphic $A_{4}$ modular flavor symmetry of level $N=3$. Within this model framework, the light neutrino masses are generated via the popular type-I seesaw mechanism in which the structure of the Dirac neutrino Yukawa couplings is decided by nonholomorphic modular forms. Unlike conventional flavor models with ad hoc flavon fields, the structure of Dirac and Majorana mass matrices is entirely determined by a modulus parameter $\tau$. We construct the predictive mass matrices for charged leptons, Dirac neutrinos, and right-handed Majorana neutrinos and show the compatibility with neutrino oscillation data by an appropriate choice of input model parameters. We find that our analysis of neutrino masses and mixing gives excellent agreement with current neutrino oscillation observables by taking normal hierarchical pattern of light neutrinos. We present numerical analysis of two sets of benchmark points explaining neutrino masses while generating the correct amount of baryon asymmetry via thermal leptogenesis. We estimate numerically the values of CP-asymmetry and examine the evolution of the lepton asymmetry by studying Boltzman equations by considering both strong and washout regimes with CP-asymmetry parameter in the range $|\varepsilon_{1}|\sim 10^{-4}$$10^{-8}$. The model predicts an effective Majorana mass in the few meV range, below current experimental bounds but within reach of next-generation $0\nu\beta\beta$ searches. The key feature of non-holomorphic $A_{4}$ modular symmetry naturally accommodates non-zero neutrino masses and mixings, minimizes the Yukawa arbitrariness, and establishes a direct connection between high-scale leptogenesis with low-energy neutrino observable parameters, thereby the model provides a testable link between neutrino flavor physics and cosmology.
@@ 1 Introduction
Even after these spectacular findings, there are some fundamental questions at the the theoretical and experimental level that remained to be answered in order to understand neutrino physics. The well-determined values of neutrino mass-square differences and mixing angles motivated to propose a simpler extension of the SM to understand the origin and structure of neutrino masses and mixing. One of the elegant idea to explain non-zero neutrino masses and mixing is popularly known as seesaw mechanism, where heavy messenger degree of freedom couple to SM neutrino, can generate small masses for light active neutrinos naturally through a suppression factor with a large messenger mass scale. The most minimal and well-studied realization is the type-I seesaw mechanism [110, 116, 150, 76], which introduces three SM-singlet right-handed neutrinos, allowing the light neutrino masses to emerge as the low-energy remnants of a high-scale Majorana mass matrix. Alternative seesaw mechanisms, including type-II [106, 146, 47, 96, 117] and type-III [72, 79], introduce scalar triplets and fermion triplets respectively, offering different model-building avenues. In one of our recent works [120], we have introduced double-seesaw mechanism to study the neutrino phenomenology.
@@
The structure of the paper is as follows. We present the adequate theoretical formulation of modular $A_{4}$ symmetry including transformation properties of particles and Yukawa couplings and assignment of proper modular weight in Sec-2, In Sec-3, we discuss the methodology of detailed numerical analysis and parameter scanning. We show the phenomenological results for neutrino oscillation observables in Sect.4 and we discuss the role of modular $A_{4}$ symmetry in estimating CP-asymmetry and study the evolution of lepton asymmetry via the Boltzmann equations to explain correct baryon asymmetry of the universe via themral leptogensis. Finally, we conclude the summary of our findings in Sect.5.
@@ 3.1 Structure of the charged lepton mass matrix
where $Y^{(0)}_{3}$ denotes the $A_{4}$ modular triplet constructed from polyharmonic Maass forms (PHMFs) of weight zero. The need for flavon fields are replaced by these modular forms and contribute to the role of effective Yukawa couplings. Their explicit Fourier expansions are provided in Eqns (A.2), (A.2).
@@ 4 Numerical Analysis of Neutrino Masses and Mixings
To begin the numerical analysis, we chose the values of ${\rm Re}[\tau]$ and ${\rm Im}[\tau]$ over a range of $|{\rm Re}[\tau]|\in$[0, 0.5] and ${\rm Im}[\tau]\in$ [0.9, 1.2] such that it satisfies the constraints of the most generic fundamental domain i.e, $|{\rm Re}[\tau]|\leq 1/2$, ${\rm Im}[\tau]>0$ and $|\tau|\geq 1$. The Yukawa couplings used in our model are evaluated using the values of ${\rm Re}[\tau]$ and Im($\tau$) from their q-expansions as given in Eqns (A.2), (A.2). We also fix phenomenologically motivated other model parameters like $|\alpha_{D},|\;|\beta_{D}|,\;|\gamma_{D}|$ over a range of [0.1, 0.99] and $\Lambda_{1},\;\Lambda_{2}$ for a range of $\Lambda_{1}\in[1.0\times 10^{13},9\times 10^{13}]$ GeV and $\Lambda_{2}\;\in\;[1.0\times 10^{14},9\times 10^{14}]$ GeV, respectively. These input values are then used to find the numerical values of $M_{D}$, $M_{R}$ and eigenvalues of the neutrino mass matrix (Eqn. 3.4).
@@
Table 4: Representative benchmark points of parameters obtained from our model with different choice of derived $M_{D}$ and fixed value of $M_{R}$. The parameters used for numerical analysis includes BP1 (2nd column) with Low $M_{D}$, $\Lambda_{1}\approx\mathcal{O}(10^{12}\;\mbox{GeV})\;\text{and}\;\;\Lambda_{2}\approx\mathcal{O}(10^{13}\;\mbox{GeV})$ and BP2 (3rd column) with High $M_{D}$, $\Lambda_{1}\approx\mathcal{O}(10^{13}\;\mbox{GeV})\;\text{and}\;\;\Lambda_{2}\approx\mathcal{O}(10^{14}\;\mbox{GeV})$.
4.1 Neutrino Observables and benchmark points
@@
$m_{1}$ (eV) $m_{2}$ (eV) $m_{3}$ (eV) $\Delta m^{2}_{21}/10^{-5}(\text{eV}^{2})$ $\Delta m^{2}_{31}/10^{-3}(\text{eV}^{2})$
$\{0.0014,0.0016\}$ $\{0.0087,0.0091\}$ $\{0.0496,0.0509\}$ $\{7.37,8.05\}$ $\{2.463,\;2.590\}$
$\delta_{CP}(^{\circ})$ $Re(\tau)$ $Im(\tau)$ $\alpha(^{\circ})$ $\beta(^{\circ})$
$\{0.00002,6.28\}$ $\{0.21,0.219\}$ $\{0.995,0.997\}$ $\{0.0037,\;360\}$ $\{0.000049,\;360\}$
$\Lambda_{1}/10^{21}$ (GeV) $\Lambda_{2}/10^{22}$ (GeV) $|\alpha_{D}|$ $|\beta_{D}|$ $|\gamma_{D}|$
$\{3.7,3.7002\}$ $\{2.353,2.355\}$ $\{0.003,0.5\}$ $\{0.006,0.609\}$ $\{0.003,0.553\}$
$|\alpha_{l}/\beta_{l}|$ $|\beta_{l}/\gamma_{l}|$
$\{0.866,0.880\}$ $\{1.65,1.70\}$
Table 5: Ranges of the various model parameters for $M_{D}\sim\mathcal{O}(1\;GeV)$ (i.e., benchmark point 1) as deduced from our Type-I $A_{4}$ model induced with modular symmetry.
$m_{1}$ (eV) $m_{2}$ (eV) $m_{3}$ (eV) $\Delta m^{2}_{21}/10^{-5}(\text{eV}^{2})$ $\Delta m^{2}_{31}/10^{-3}(\text{eV}^{2})$
$\{0.0014,0.0017\}$ $\{0.008,0.009\}$ $\{0.049,0.0510\}$ $\{6.99,8.05\}$ $\{2.463,\;2.606\}$
$\delta_{CP}(^{\circ})$ $Re(\tau)$ $Im(\tau)$ $\alpha(^{\circ})$ $\beta(^{\circ})$
$\{96.0,422.0\}$ $\{0.21,0.22\}$ $\{0.99,0.999\}$ $\{0.0037,360\}$ $\{0.00005,360\}$
$\Lambda_{1}/10^{22}$ (GeV) $\Lambda_{2}/10^{23}$ (GeV) $|\alpha_{D}|$ $|\beta_{D}|$ $|\gamma_{D}|$
$\{8.68,\;8.7\}$ $\{6.6,\;6.67\}$ $\{0.006,0.534\}$ $\{0.005,0.778\}$ $\{0.005,0.669\}$
$|\alpha_{l}/\beta_{l}|$ $|\beta_{l}/\gamma_{l}|$
$\{1.19,\;1.23\}$ $\{0.72,\;0.75\}$
Table 6: Ranges of the various model parameters for $M_{D}\sim\mathcal{O}(10\;GeV)$ (i.e., benchmark point 2) as deduced from our Type-I $A_{4}$ model induced with modular symmetry.
@@
[…]
@@ 4.2 Correlation plots and modular dependence
On the other hand, Fig. 2 depicts the correlation between the real and imaginary parts of the modulus parameter $\tau$. Higher order of $M_{D}$ masses indicate the enhanced modular space between the real and imaginary planes of the modulus parameter that transforms under the modularity condition in the upper half plane $\mathcal{H}$. This gives us the idea about the sensitivity of $M_{D}$ towards the modular space as derived from our model. Fig. 3 illustrates the robust correlation between the square of the sine of solar and reactor mixing angles within the allowed $3\sigma$ region. This correlation is crucial for the viability of our model as it impacts the parameter space of the other neutrino observables and model derived parameters. In Fig. 4, we have shown the strong correlation between the square of the sine of the solar and atmospheric mixing angles as it influences the neutrino oscillation parameters.
In Fig. 5, we illustrate the correlation of the first two light neutrino mass eigenvalues, $m_{1}$ and $m_{2}$ such that it satisfies $\Delta m^{2}_{12}/10^{-5}\;\in\;[6.92,\;8.05]$ as given in NuFIT 6.0. In the first case BP1 ($M_{D}\sim\mathcal{O}(1\;GeV)$), the parameter space is narrower and the points are more clustered, while for BP 2 ($M_{D}\sim\mathcal{O}(10\;GeV)$), the parameter space has a wider space, indicating how the hierarchical structures are induced by the mass order of Dirac neutrino masses. Thus we can get a more predictive correlation space for higher $M_{D}$. Similarly, in Fig. 6, the correlation between the lightest and the heaviest masses is shown for the normal hierarchy. It is to be noted that our model predicts the hierarchy of the light neutrino masses as $m_{1}<m_{2}<m_{3}$. Thus, the discrete and scattered points in Figs. 5 and 6 reflect how $m_{1}$, $m_{2}$ and $m_{3}$ are highly constrained after the implementation of the type-I seesaw model with modular $A_{4}$ symmetry for both BP1 and BP2.
In Fig. 7, we find that the solar and atmospheric mass squared differences as obtained from our model show a very strong correlation indicating that the parameter values of $\Delta m^{2}_{12}$ and $\Delta m^{2}_{23}$ derived from our type-I seesaw model agree well its the experimental values. This hints us that our model is highly predictive and testable for ongoing and upcoming neutrino experiments. In Fig. 8, we have shown the correlation between the solar mixing angle $\sin^{2}\theta_{12}$ and the solar mass squared difference $\Delta m^{2}_{12}$.
[…]
For both the above mentioned benchmark points BP 1 and BP 2, the predicted value of the lightest neutrino mass is of the order of $\mathcal{O}(\text{meV})$ and the element of the effective neutrino mass $|m_{ee}|$ is predicted to be small (around a few meV), which is well below the current experimental bounds but potentially testable by next-generation and beyond ton-scale experiments. In both cases, our predicted value of $\delta_{CP}$ falls within the range of parameter parameters of $3\sigma$. Also, it should be noted that no data points were generated for the inverted hierarchy, which satisfies the constraints of mass-squared differences.
[…]
@@
Experiment Isotope $T^{0\nu}_{1/2}\;[10^{25}\;\text{yrs}]$ $m_{ee}\;(\text{meV})$ Ref
GERDA+MJD+L200 $^{76}\text{Ge}$ 1919 7520075-200 [12]
KamLAND-Zen 800 $^{136}\text{Xe}$ $38$ 2812228-122 [7]
CUORE $^{130}\text{Te}$ $3.8$ 7024070-240 [14]
@@
$T^{0\nu}_{1/2}$ $m_{ee}\;(\text{meV})$
CUPID $\text{Li}_{2}^{100}\text{MoO}_{4}\;\text{(crystal)}$ $1.1\times 10^{27}\;\text{yrs}$ 122012-20 [22]
LEGEND-1000 $^{76}\text{Ge}$ $1.3\times 10^{28}\;\text{yrs}$ 9219-21 [9]
nEXO $^{136}\text{Xe}$ $7.4\times 10^{27}\;\text{yrs}$ 6276-27 [15]
Future Beyond Ton-Scale Experiments of $0\nu\beta\beta$ decay
@@
$T^{0\nu}_{1/2}$ $m_{ee}\;({\rm meV})$
LEGEND-6000 $^{76}\text{Ge}$ $10^{29}\;\text{ yrs}$ 363-6 [1]
ORIGIN-X $^{136}\text{Xe}$ $10^{30}\;\text{ yrs}$ <2<2 [80]
THEIA $^{130}\text{Te}(^{136}\text{Xe})$ $1.1(2.0)\times 10^{28}\;\text{yrs}$ <6.3(5.6)<6.3(5.6) [84]
$90\%\;\text{C.L.}\;\text{for }\;\text{Te (Xe)}$